OpenStudy (anonymous):

A rectangle has a perimeter of 80cm. If its width is x, express its length and area in terms of x. Find also the maximum area.

OpenStudy (anonymous):

P = 80 80 = X + L L = 80-X A = X*L A = X*(80-x) maximum area is found by taking the derivative of area equation, which is, X*(80-X), once derived, set it equal to 0 and solve for X